OpenStudy (anonymous):

In a dice game if you roll a 2, 4, or 6 you get the value of the die. If you roll as 1, 3, or 5 you lose $5. What is the expected value of the game?

OpenStudy (anonymous):

multiply the amount you win times the probability you win it and add it up

OpenStudy (anonymous):

the probability you get any of those outcome is \(\frac{1}{6}\) so compute via \[2\times \frac{1}{6}+4\times \frac{1}{6}+6\times \frac{1}{6}-5\times \frac{1}{6}-5\times \frac{1}{6}-5\times \frac{1}{6}\] \[=\frac{2+4+6-5-5-5}{6}\]

OpenStudy (anonymous):

i get \[\frac{12-15}{6}=\frac{-3}{6}=-\frac{1}{2}\] meaning on average you lose half a dollar every time you play

OpenStudy (unklerhaukus):

if {2,4,6} you win you can expect to win $4 as 4 is the average of {2,4,6} if{1,3,5} you loose you loose $5 since there are as many terms in each case each is al likely, so you can expect to win as much as you loose, the average of {4,-5}=(4-5)/2 =-1/2